# Reason-to-Rank: Learning to Rank through Reasoning-Based Knowledge Distillation

### v2

@@ 4.1 Datasets and Evaluation Metrics
4.2 Models
4.2.1 Teacher models
We utilize large language models (LLMs) as teacher models to generate reasoning annotations: GPT-4 Achiam et al. (2023), Claude The (claude-3-5-sonnet-20240620) and Gemini Team et al. (2023) (gemini-1.5-flash). These advanced LLMs known for their strong reasoning and language-understanding capabilities.
@@ 4.2.2 Student models
we also evaluated different reasoning strategies in the R2R student models, specifically analyzing the "w/o Reasoning" (without reasoning), "w/ direct relevance reasoning" (with direct relevance reasoning), "w/ comparison reasoning" (with comparison reasoning), and "w/ direct relevance & comparison" (with both direct relevance and comparison reasoning) variants.
As shown in Table  1, models without reasoning (using pairwise or listwise loss) perform at a baseline level, with NDCG@5 scores of 73.2% and 73.8 on DL19 and 70.8% and 71.0% on DL20. Adding reasoning significantly boosts performance. For example, direct relevance reasoning improves NDCG@5 to 74.5% on DL19 and 70.1% on DL20 by providing direct relevance relevance explanations. Comparison reasoning, which assesses the relative importance between documents, raises NDCG@5 to 74.1% on DL19 and 72.3% on DL20. Combining both types of reasoning yields the best results, with NDCG@5 scores of 75.4% on DL19 and 72.4% on DL20.
@@
4.2.3 Comparison with Baselines
Our student models, trained with direct relevance and comparison reasoning, achieve competitive performance compared to the previous state-of-the-art student models. As shown in Table  1,the improvements over the baseline models are statistically significant, with p-values less than 0.05.
@@ 4.2.3 Comparison with Baselines
RankGPT DeBERTa Sun et al. (2023): A transformer model using disentangled attention to better capture long-range dependencies, improving reranking by understanding query-document context.
MonoT5 Nogueira et al. (2020): A T5-based model treating reranking as a sequence generation task, leveraging its encoder-decoder architecture for better relevance scoring.
RankVicuna, RankZephyr and APEER Pradeep et al. (2023a, b); Jin et al. (2024): RankVicuna is a distilled version of RankGPT, designed for zero-shot listwise reranking with fewer parameters, balancing efficiency and performance. RankZephyr a state-of-the-art zero-shot listwise reranking model that adapts RankGPT prompts for robust performance with minimal task-specific training. APEER share the same structure with RankZephyr but have an automated prompt generation for large language models.
4.2.4 R2R Student Model vs. Baselines
@@
4.2.5 Performance on Reasoning-Intensive Tasks (BRIGHT Dataset)
The BRIGHT dataset Su et al. (2024) is specifically designed to evaluate models on reasoning-intensive retrieval tasks, simulating real-world scenarios where complex reasoning is required to determine document relevance. We report the Inst-LSu et al. (2024, 2022) model as a baseline because it is fine-tuned with task-specific instructions, making it particularly well-suited for tasks requiring deeper reasoning. Table 2 shows the NDCG@5 results on the BRIGHT dataset. In the Earth Science domain, our R2R model achieves an NDCG@5 of 31.0%, closely approaching the teacher model’s 37.1%. Similarly, in Psychology, the student model attains 28.9%, compared to the teacher’s 34.2% and BM25’s 11.6%. However, we observe that in certain domains, such as Pony and AoPS, the performance of our student and teacher models is relatively low.
Domain BM25 Inst-L Su et al. (2024) R2R GPT-4 Teacher
@@ 4.2.5 Performance on Reasoning-Intensive Tasks (BRIGHT Dataset)
Table 2: NDCG@5 results on the BRIGHT dataset.
5 Ablation Study
5.1 Quality of Generated Reasons
We assess the quality of the generated reasons using BLEU and ROUGE-L as presented in Table 3 compared with the teacher model’s reason. The baseline model here represents the model trained without the rank loss function and has the same input as ours. On the DL19 dataset, the student model achieves a BLEU score of 21.4, compared to the baseline’s 18.0, and a ROUGE-L score of 36.8 versus 32.5. Similar improvements are observed on DL20, where the student model attains BLEU and ROUGE-L scores of 24.0 and 38.2, respectively, surpassing the baseline’s 19.5 and 34.0.
@@ 5.1 Quality of Generated Reasons
Table 3: Evaluation of reason generation using BLEU and ROUGE-L scores for student models trained with and without the ranking loss function. The ’Student Model w/o Rank Loss’ is trained without the rank loss function and has the same input as the ’Student Model w/ Rank Loss’.
5.2 Impact of Reasoning Strategies
In Table 4, by removing comparison reasoning results in a performance drop on DL19 from 75.3 to 74.1 (NDCG@5). Similarly, removing direct relevance reasoning causes a performance decrease on DL20 from 72.3 to 72.2 (NDCG@5). Both reasoning strategies contribute to performance improvements.
@@ 5.2 Impact of Reasoning Strategies
Table 4: Ablation study: Impact of direct relevance and comparison reasoning on NDCG@5/@10 for DL19 and DL20.
5.3 Effects of Training Data Size
@@
Figure 4 i shows the performance of small models trained on 100, 1,000, and 2,000 examples. Contrary to expectations, increasing data size from 1,000 to 2,000 does not consistently improve performance and sometimes slightly reduces it. We find similar trends across various small models, including Mistral and LLaMA variants, highlighting that adding more data does not always guarantee better performance.
@@
Figure 4: Student model performance for different training data sizes on DL19 and DL20.
BM25 50.6 48.0 41.0 59.5 44.2 39.5 30.8 40.7 31.8
Rank GPT-4 Model
Rank GPT-4 (w/o reasoning) 75.6 70.6 54.0 85.5 38.5 57.6 36.6 57.6 47.1
+ pairwise reasoning 75.3 69.9 52.9 83.3 38.9 55.9 36.1 52.9 47.0
+ comparison reasoning 76.3 71.1 53.4 84.6 38.2 56.5 36.5 56.5 48.2
Claude Model
Claude (w/o reasoning) 72.0 68.8 50.8 82.8 36.1 54.4 36.4 50.8 42.5
+ pairwise reasoning 70.6 68.2 50.3 82.2 35.9 55.4 35.6 52.4 42.5
+ comparison reasoning 71.3 69.8 51.1 83.3 36.4 55.1 36.7 51.5 43.2
Gemini Model
Gemini (w/o reasoning) 71.1 68.5 50.1 83.0 35.6 53.7 35.5 50.1 42.9
+ pairwise reasoning 69.1 67.0 51.1 81.6 34.8 52.9 36.9 51.3 43.9
+ comparison reasoning 70.2 68.2 50.2 82.4 35.9 53.4 35.4 51.1 42.9
Table 5: NDCG@5 performance (in percentage) for different models with various reasoning prompts across multiple datasets. Bold numbers indicate the best performance within each group of models. indicates statistically significant improvement over the base model without reasoning prompts (p $<$ 0.05). indicates statistically significant improvement over both the base model and the single reasoning prompt variants (p $<$ 0.05).
6 Discussion
Our experiments show that the R2R framework effectively enhances document reranking by incorporating direct relevance and comparative reasoning. The student models achieve competitive results while significantly reducing computational resources compared to large teacher models. The statistically significant improvements confirm that reasoning strategies effectively enhance model performance. An important aspect of R2R is its ability to not only rank documents but also generate coherent and relevant reasoning for its rankings. This dual capability suggests that the model could be applied to other tasks requiring both ranking and reasoning generation.
While our current work centers on document reranking in information retrieval, the integration of reasoning with ranking decisions holds promise for broader applications. For instance, in recommendation systems, the model could rank products or services while providing explanations for its choices, enhancing user trust. Similarly, in question answering or summarization tasks, generating reasoning alongside answers or summaries could improve understandability Jiang et al. (2024). However, future work is needed to explore the model’s performance in these varied contexts and verify its generalizability beyond document reranking McElfresh et al. (2022); Rahmani et al. (2022).
7 Conclusion
In this paper, we introduced ReasoningRank, a novel approach to document reranking that incorporates direct relevance and comparative reasoning. Our method leverages the power of large language models to generate interpretable reasons, which are then distilled into smaller, more efficient student models. Our experiments show the effectiveness of this approach, achieving high performance across multiple datasets while maintaining interpretability.
Limitations
While our approach shows strong performance, it has several limitations. First, the reliance on large language models for generating reasoning limits the applicability of our method in resource-constrained environments. Second, the quality of the generated reasons heavily depends on the quality of the prompts, which requires careful crafting. Future research should address these limitations by exploring more efficient models and automated prompt-generation techniques.
Ethics Statement
This work enhances the transparency and accessibility of reasoning capabilities in AI by distilling knowledge from LLMs to smaller models. While this improves efficiency, we acknowledge the ethical concerns of bias in LLMs and the need for careful mitigation in real-world applications.
Appendix A Student Training Parameters
@@
Table 6: Training configuration hyperparameters.
Appendix B Teacher Model Parameters
Hyperparameter Value
Model GPT
Temperature 1.0
Top p 0.9
Table 7: Teacher model hyperparameter settings.
Appendix C NDCG@5 for Different Models
Table 8: NDCG@5 for different models and training data sizes on various datasets.
Model Train Size DL 19 DL 20
Mistral v0.1 100 69.69 76.33
1000 71.03 77.06
2000 70.63 73.99
Mistral v0.3 100 69.26 79.55
1000 71.56 79.14
2000 69.86 76.96
LLaMA 3.1 8B 100 69.63 75.85
1000 72.87 77.03
2000 73.37 77.60
LLaMA 3 8B 100 69.57 77.15
1000 73.76 75.35
2000 71.75 75.55
LLaMA 2 7B 100 71.50 79.89
1000 72.19 72.95
2000 72.18 75.36
Appendix D Model Behavior Analysis
In the reasoning generation tasks, the model’s performance varied across direct relevance reasoning, comparison reasoning, and their combination. Below, we provide an analysis of unexpected behavior based on 2,000 queries.
Repetition.
The repetition metric measures the occurrence of duplicate passage identifiers. Across the tasks, we find:
  • Pairwise reasoning: 63 duplicate passage identifiers (3%).
  • Comparison reasoning: 36 duplicate passage identifiers (1.5%).
  • Combined reasoning: 47 duplicate passage identifiers (2%).
Missing Documents.
Missing documents were most frequent in direct relevance reasoning:
  • Pairwise reasoning: 642 missing documents (32% of queries).
  • Comparison reasoning: 227 missing documents (11% of queries).
  • Combined reasoning: 328 missing documents (16% of queries).
Appendix E API Cost Analysis
Utilizing large language models (LLMs) like GPT-4 as teacher models involves significant computational and monetary costs. In this section, we analyze these costs based on our experiments.
We calculated the average number of tokens, the number of API requests, and the estimated cost per query for different operations involving the teacher model. The costs are based on the pricing provided by the API service at the time of our experiments 9.
@@ Appendix E API Cost Analysis
Table 9: Average token usage and cost per query using GPT-4 for different reasoning tasks.
Appendix F Prompts for Reasoning
Table 10: Basic prompt, direct relevance reasoning prompt, and comparison reasoning prompt.
Basic Prompt:
I will provide you with num passages, each indicated by a number identifier [].
For each passage, briefly generate your reasoning process as follows:
1. Judge whether the passage is not applicable (not very common), where the query does not meet the
premise of the passage.
2. Check if the query contains direct evidence. If so, judge whether the query meets or does not meet
the passage.
3. If there is no direct evidence, try to infer from existing evidence and answer one question:
If the passage is ranked in this order, is it possible that a good passage will miss such information?
If impossible, then you can assume that the passage should not be ranked in that order.
Otherwise, it should be ranked in that order.
Direct relevance Reasoning Prompt:
Then, read every passage one by one and find the sentence where the method is the direct answer
for the query and output the sentence. Then, rank the passages based on their relevance to the query.
Give highest priority to passages that provide a clear and direct definition or explanation of the
keywords related to the query.
Consider both the detailed information and any relevant background context provided in each passage.
Provide clear and concise reasons for the ranking, highlighting the specific parts of the passages that
influenced your decision.
Make sure to ignore any irrelevant information and focus on content directly related to the query.
In addition to direct reasons for ranking each passage, also consider listwise reasons.
Comparison Reasoning Prompt:
A listwise reason involves comparing passages with each other to determine their relative importance.
Specifically, compare passages to identify which ones provide more comprehensive, relevant, and accurate
information in relation to the query.
When providing listwise reasons, mention specific comparative insights, such as why one passage might be
more relevant than another based on the overall context and detail provided.
Return Type:
Search Query: query. Rank the num passages above based on their relevance to the search query and
the extracted keywords. The passages should be listed in descending order using identifiers.
The most relevant passages should be listed first. The output should be in JSON format.
For each passage, provide the extracted keywords and detailed reasons for its ranking.
Ensure to separate direct reasons and listwise reasons clearly.
Mention specific parts of the passage that influenced your decision.
Ensure the reasons are clear, concise, and directly related to the query,
balancing direct definitions or explanations and relevant background information.
Only output the JSON structured format.

### v4

@@ 4.1. Datasets and Evaluation Metrics
4.2. Models
4.2.1. Teacher models
We utilize large language models (LLMs) as teacher models to generate reasoning annotations: GPT-4 (Achiam et al., 2023), Claude (The, [n. d.]) (claude-3-5-sonnet-20240620) and Gemini (Team et al., 2023) (gemini-1.5-flash). These advanced LLMs known for their strong reasoning and language-understanding capabilities.
@@ 4.2.2. Student models
we also evaluated different reasoning strategies in the R2R student models, specifically analyzing the ”w/o Reasoning” (without reasoning), ”w/ direct relevance reasoning” (with direct relevance reasoning), ”w/ comparison reasoning” (with comparison reasoning), and ”w/ direct relevance & comparison” (with both direct relevance and comparison reasoning) variants.
As shown in Table  2, models without reasoning (using pairwise or listwise loss) perform at a baseline level, with NDCG@5 scores of 73.2% and 73.8% on DL19 and 70.8% and 71.0% on DL20. Adding reasoning significantly boosts performance. For example, direct relevance reasoning improves NDCG@5 to 74.5% on DL19 and 70.1% on DL20 by providing direct relevance relevance explanations. Comparison reasoning, which assesses the relative importance between documents, raises NDCG@5 to 74.1% on DL19 and 72.3% on DL20. Combining both types of reasoning yields the best results, with NDCG@5 scores of 75.4% on DL19 and 72.4% on DL20.
@@
4.2.3. Comparison with Baselines
Our student models, trained with direct relevance and comparison reasoning, achieve competitive performance compared to the previous state-of-the-art student models. As shown in Table  2,the improvements over the baseline models are statistically significant, with p-values less than 0.05.
@@ 4.2.3. Comparison with Baselines
RankGPT DeBERTa (Sun et al., 2023): A transformer model using disentangled attention to better capture long-range dependencies, improving reranking by understanding query-document context.
MonoT5 (Nogueira et al., 2020): A T5-based model treating reranking as a sequence generation task, leveraging its encoder-decoder architecture for better relevance scoring.
RankVicuna, RankZephyr and APEER (Pradeep et al., 2023a, b; Jin et al., 2024): RankVicuna is a distilled version of RankGPT, designed for zero-shot listwise reranking with fewer parameters, balancing efficiency and performance. RankZephyr a state-of-the-art zero-shot listwise reranking model that adapts RankGPT prompts for robust performance with minimal task-specific training. APEER share the same structure with RankZephyr but have an automated prompt generation for large language models.
4.2.4. R2R Student Model vs. Baselines
@@
For Touche, it is shown that it is a argument retrieval or debate scenario dataset, and BM25 tends to perform surprisingly well in such “keyword-matching” strong correlation scenarios; furthermore, some models (e.g., RankVicuna) may be better adapted in argument mining Furthermore, some models (e.g., RankVicuna) may be better adapted for argument mining. For Robust04, this means that it covers a wider range of topics and outdated documents, where RankZephyr’s multi-iteration listwise training is more advantageous.
4.2.5. Performance on Reasoning-Intensive Tasks (BRIGHT Dataset)
The BRIGHT dataset (Su et al., 2024) is specifically designed to evaluate models on reasoning-intensive retrieval tasks, simulating real-world scenarios where complex reasoning is required to determine document relevance. We report the Inst-L(Su et al., 2024, 2022) model as a baseline because it is fine-tuned with task-specific instructions, making it particularly well-suited for tasks requiring deeper reasoning. Table 3 shows the NDCG@5 results on the BRIGHT dataset. In the Earth Science domain, our R2R model achieves an NDCG@5 of 31.0%, closely approaching the teacher model’s 37.1%. Similarly, in Psychology, the student model attains 28.9%, compared to the teacher’s 34.2% and BM25’s 11.6%. However, we observe that in certain domains, such as Pony and AoPS, the performance of our student and teacher models is relatively low.
Table 3. NDCG@5 results on the BRIGHT dataset.
Domain BM25 Inst-L(Su et al., 2024) R2R GPT-4 Teacher
@@ 4.2.5. Performance on Reasoning-Intensive Tasks (BRIGHT Dataset)
5. Ablation Study
5.1. Quality of Generated Reasons
We assess the quality of the generated reasons using BLEU and ROUGE-L as presented in Table 4 compared with the teacher model’s reason. The baseline model here represents the model trained without the rank loss function and has the same input as ours. On the DL19 dataset, the student model achieves a BLEU score of 21.4%, compared to the baseline’s 18.0%, and a ROUGE-L score of 36.8% versus 32.5%. Similar improvements are observed on DL20, where the student model attains BLEU and ROUGE-L scores of 24.0% and 38.2%, respectively, surpassing the baseline’s 19.5% and 34.0%.
Table 4. Evaluation of reason generation using BLEU and ROUGE-L scores for student models trained with and without the ranking loss function. The ’Student Model w/o Rank Loss’ is trained without the rank loss function and has the same input as the ’Student Model w/ Rank Loss’.
@@ 5.1. Quality of Generated Reasons
5.2. Impact of Reasoning Strategies
In Table 5, by removing comparison reasoning results in a performance drop on DL19 from 75.3% to 74.1% (NDCG@5). Similarly, removing direct relevance reasoning causes a performance decrease on DL20 from 72.3% to 72.2% (NDCG@5). Both reasoning strategies contribute to performance improvements.
Table 5. Ablation study: Impact of direct relevance and comparison reasoning on NDCG@5/@10 for DL19 and DL20.
@@ 5.2. Impact of Reasoning Strategies
5.3. Effects of Training Data Size
@@
Figure 5 shows the performance of small models trained on 100, 1,000, and 2,000 examples. Contrary to expectations, increasing data size from 1,000 to 2,000 does not consistently improve performance and sometimes slightly reduces it. Similar trends across various small models, including Mistral and LLaMA variants, highlight that adding more data does not always guarantee better performance.
@@
Figure 5. Student model performance for different training data sizes on DL19 and DL20.
Table 6. Impact of different $(\alpha,\beta,\gamma)$ combinations on model performance. We report the NDCG (averaged at a certain cutoff), BLEU, and ROUGE-L scores under different weighting strategies. The best performance is in bold.
Hyperparameters Performance
$\alpha$ $\beta$ $\gamma$ NDCG BLEU ROUGE-L
0.2 0.2 0.6 0.78 0.74 0.72
0.3 0.3 0.5 0.82 0.76 0.75
0.4 0.4 0.2 0.86 0.79 0.78
0.5 0.5 0.1 0.84 0.77 0.76
0.6 0.6 0.05 0.80 0.73 0.74
5.4. Hyperparameter Tuning.
We conducted a small-scale grid search over $\alpha$, $\beta$,$\gamma$ on a held-out validation set. Table 6 shows the results for different weighting strategies. We find that ($\alpha$=0.4,$\beta$=0.4,$\gamma$=0.2) yields the best trade-off among NDCG, BLEU, and ROUGE-L, and we thus adopt these settings for all subsequent experiments.
5.5. Case Study
In addition to ranking improvements, the generated explanations offer interpretability benefits. For example, in several case studies, see Table 7, we observed that the comparison reasoning often highlights subtle differences between top-ranked documents, which can assist developers in diagnosing ranking errors. Although our current evaluation of explanations relies on automated metrics (BLEU, ROUGE), a small-scale human evaluation is planned in future work to quantitatively assess readability, consistency, and practical utility of these explanations.
Table 7. An illustrative case study showing how direct and comparison reasoning enhance interpretability for a medical-related query. In this example, Document A is ultimately deemed more relevant due to its detailed coverage of insulin’s regulatory mechanism, whereas Document B offers a more general discussion of hormones.
Query: ”How does insulin regulate blood sugar?”
Document A (snippet)
“Insulin is produced by the pancreas and helps cells take in glucose from the bloodstream. By binding to insulin receptors, it initiates a signaling cascade that lowers blood glucose levels…”
Direct Reasoning:
“This document directly addresses how insulin facilitates cellular glucose uptake, which aligns precisely with the question on regulating blood sugar.”
Comparison Reasoning vs. Document B:
“Compared to Document B, Document A provides more explicit details on insulin’s binding mechanism and the biochemical pathway, making it more relevant for understanding exactly how insulin regulates glucose.”
Document B (snippet)
“Hormones like insulin and glucagon maintain metabolic balance. Insulin generally lowers blood sugar levels, while glucagon raises them, ensuring homeostasis…”
Direct Reasoning:
“This document briefly mentions insulin’s role in lowering blood sugar, but primarily focuses on the broader hormone interplay.”
Comparison Reasoning vs. Document A:
“Although Document B is relevant, it emphasizes multiple hormones and offers less specific detail on the mechanism of insulin. Document A is therefore more closely aligned with how insulin regulates blood sugar.”
6. Discussion and Conclusion
Our experimental results demonstrate that Reason-to-Rank (R2R) effectively enhances document reranking by integrating both direct relevance and comparative reasoning. Compared to relying solely on a large teacher model (e.g., GPT-4), our distilled student models offer competitive performance while drastically reducing computational cost, making the framework suitable for real-world, high-throughput settings. The statistically significant gains confirm that incorporating explicit reasoning structures—both pointwise and pairwise—meaningfully contribute to improved ranking accuracy.
A natural concern is whether the observed performance improvements stem merely from generating longer output text rather than the intrinsic value of direct and comparison reasoning. We argue that the benefit of our approach lies in the structured content of the explanations: by explicitly contrasting documents (comparison reasoning) and detailing query-document alignment (direct reasoning), the model captures fine-grained distinctions that standard ranking scores might overlook. Although we have not conducted a controlled experiment to equalize text length, preliminary analyses indicate that even when constraining the output to a fixed token budget, models incorporating both reasoning types consistently outperform those without. Future work will include a more rigorous ablation study and a human evaluation to further assess the usefulness of the generated rationales in aiding system diagnosis and debugging.
Notably, the proposed student models not only output document rankings but also produce coherent and relevant justifications for each document’s position. This dual capacity to rank and explain suggests broader applicability beyond traditional ad-hoc retrieval. For example, in recommendation systems, R2R could prioritize items while elucidating the rationale (e.g., product features or user context), thereby increasing user trust and satisfaction. Similarly, in question answering or summarization tasks, generating reasoning alongside answers may boost transparency and user acceptance (Jiang et al., 2024). However, systematically evaluating these broader domains remains as future work to verify the generalizability of our approach (McElfresh et al., 2022; Rahmani et al., 2022).
Overall, the synergy between direct relevance and comparative reasoning leads to both practical efficiency and rich interpretability. Our findings point to an emerging trend in IR and NLP systems where explicit reasoning improves performance metrics and contributes to more explainable and user-friendly outcomes.
In this paper, we introduced Reason-to-Rank (R2R), a novel framework that unifies direct relevance and comparison reasoning for document reranking. We leveraged large language models (LLMs) as teacher models to generate high-quality explanations and rankings, then distilled both the ranking expertise and textual justifications into smaller student models. Experimental results across multiple benchmarks illustrate that R2R not only achieves strong reranking effectiveness but also delivers interpretable rationales for why one document is ranked above another. Our study underscores the value of distilling explicit reasoning strategies from large LLMs and highlights the benefits of merging pointwise and pairwise perspectives.
7. Limitations and Future Work
Despite the encouraging results, there are several limitations to consider: Dependence on large LLMs for teacher data. Our approach relies on powerful teacher models (e.g., GPT-4) to generate high-quality explanations. Obtaining robust supervision signals could be challenging in resource-constrained environments or specialized domains where such models are not readily available. Also, using GPT-4 can be expensive (approx.\$0.2 per query with 4,000+ tokens). While knowledge distillation substantially reduces cost at inference time, the distilled models are still large and thus not trivial for real-time production in specific scenarios. Prompt quality and model biases. The effectiveness of the distilled reasoning hinges on carefully designed prompts. If prompts fail to capture key aspects of the query or document, the generated explanations might be shallow or incorrect. Additionally, teacher models can introduce biases, which might propagate into student models during distillation.
Broader applicability beyond reranking. While we have discussed possible extensions to recommendation, QA, or summarization tasks, the actual performance and adaptation strategy in these domains remain largely unexplored. Thorough empirical studies are required to confirm that the proposed dual reasoning approach generalizes effectively. Lack of Human Evaluation for Explanations We only measured textual explanations with BLEU/ROUGE against the teacher outputs, which themselves are automatically generated. This does not guarantee that the explanations are genuinely helpful or correct from a human perspective. Generalization beyond Document Ranking. Though we focus on ad-hoc retrieval, the idea of distilling direct and comparative reasoning could benefit other tasks like question answering or recommender systems. More thorough domain-specific experiments are necessary to confirm the general applicability of our approach.Confounding Factors of Explanation Length.While our ablation controls for token length, further studies are needed to rigorously isolate how much of the ranking improvement stems from the reasoning structure vs. extended context. Understanding how best to prompt for concise but effective rationales remains an open question.
In future work, we plan to investigate more efficient teacher–student paradigms (e.g., smaller teacher models, automated prompt generation) and conduct domain-specific evaluations to further validate the versatility of R2R. We believe these efforts can refine the balance between interpretability and computational cost, extending the framework’s applicability across diverse IR and NLP scenarios.
8. Appendix
8.1. Student Training Parameters
Table 8. Training configuration hyperparameters.
@@ 8.1. Student Training Parameters
8.2. API Cost Analysis
We calculated the average number of tokens, the number of API requests, and the estimated cost per query for different operations involving the teacher model, see in Table 9. The costs are based on the pricing provided by the API service at the time of our experiments.
Table 9. Average token usage and cost per query using GPT-4 for different reasoning tasks.